Need help with networking problem
Imagine you have a server containing a movie segmented into chunks and encoded at constant bitrates of
200, 500 and 1000 Kbps. With the appropriate use of diagrams and clear descriptive text:
• How many bytes make up the chunks representing both 2 seconds and 8 seconds of content at the
200, 500 and 1000 Kbps representation rates? (Six different types of chunks in total.)
• How many seconds would each of these chunks take to send if they were encapsulated in UDP/IP
frames and transmitted at precisely 1600Kbps (at the IP layer), with no other overheads and no packet
losses?
Answer #1

2 seconds -

For bit rate of 200 Kbps Since 200 Kb generated in 1 second Hence 400Kb will be produced in 2 seconds

For bit rate of 500 Kbps Since 500 Kb generated in 1 second Hence 1000Kb will be produced in 2 seconds

For bit rate of 1000 Kbps Since 1000 Kb generated in 1 second Hence 2000Kb will be produced in 2 seconds

8 seconds -

For bit rate of 200 Kbps Since 200 Kb generated in 1 second Hence 1600Kb will be produced in 8 seconds

For bit rate of 500 Kbps Since 500 Kb generated in 1 second Hence 4000Kb will be produced in 8 seconds

For bit rate of 1000 Kbps Since 1000 Kb generated in 1 second Hence 8000Kb will be produced in 8 seconds

Now consider that the chunks are transmitted at 1600Kbps hence the Chunk of X kb will require X / 1600 seconds

Chunk of size 400Kb - 400/1600 = 0.25 sec

Chunk of size 1000Kb - 1000/1600 = 0.625 sec

Chunk of size 2000Kb - 2000/1600 = 1.25 sec

Chunk of size 1600Kb - 1600/1600 = 1 sec

Chunk of size 4000Kb - 4000/1600 = 2.5 sec

Chunk of size 8000Kb - 8000/1600 = 5 sec

I have tried to explain it in very simple language and I hope that i have answered your question satisfactorily.Leave doubts in comment section if any